a procedure similar to that described in this experiment can be used to precipitate Cl- as AgCl. A sample containing chloride weighing .207g was precipitate with Ag+, and the AgCl produced weighed .472g. Calculate the mass % of Cl- in the original sample
1
Expert's answer
2017-11-13T09:09:06-0500
〖Ag〗^++〖Cl〗^-⟹AgCl↓
m(〖Cl〗^- )= (M(Cl)×m(AgCl))/(M(AgCl))= (35.5×0.472)/142.5=0.108 (g)
w(〖Cl〗^- )= m(〖Cl〗^- )/(m(sample))×100%= 0.108/0.207×100%=52.2 %
Answer: 52.2 %.